# RAS: Retrieval-And-Structuring for Knowledge-Intensive LLM Generation

### v1

@@ 5 Conclusion
We presented RAS (Retrieval-And-Structuring), a novel framework that enhances knowledge-intensive language generation through iterative retrieval and knowledge structuring. Our approach introduces key innovations including theme-scoped retrieval, action planning, dynamic knowledge structuring, and graph-augmented answering. Comprehensive experiments demonstrate RAS’s superior performance over existing approaches across diverse benchmarks. Looking ahead, RAS’s graph-based architecture enables seamless integration with external knowledge graphs and human-in-the-loop curation, making it particularly valuable for specialized domains requiring high accuracy and reliability. We discuss limitations of RAS and future work in Appendix A. Our code can be found at https://github.com/pat-jj/Retrieval-And-Structure.
Impact Statements
The RAS framework has several potential positive impacts on both research and practical applications. First, by improving knowledge-intensive language generation through structured knowledge representation and iterative reasoning, RAS could enhance the reliability and factual accuracy of AI systems in critical domains like education, research assistance, and technical documentation. The theme-scoped retrieval mechanism also offers computational efficiency benefits, potentially reducing the energy consumption and computational resources required for large-scale knowledge retrieval tasks.
However, there are important considerations regarding potential risks and challenges. The framework’s effectiveness in knowledge structuring could potentially amplify existing biases in training data and knowledge sources. Additionally, while RAS improves retrieval efficiency, it may create barriers for low-resource languages or domains where structured knowledge bases and thematic classifications are limited. These concerns necessitate careful consideration of data sources and regular evaluation of the system’s impact across different user groups and application contexts.
@@ Appendix A Limitations & Future Work
Knowledge Extraction Quality: The effectiveness of both planning and answer generation depends on the quality of text-to-triple extraction. Despite using a Flan-T5-based model for triple extraction, implementing a more sophisticated approach – such as a dedicated Named Entity Recognition and Relation Extraction (NER-RE) pipeline – could potentially enhance RAS’s reasoning capabilities and knowledge representation.
Theme Scoping Limitations: The current classifier was trained on DBPedia-298, an older dataset that may not adequately represent the thematic distribution of contemporary document collections. This temporal gap particularly affects performance when processing newer corpora, as evidenced by our experiments with Wikipedia 2020 dumps. The development of a huge, up-to-date multilabel document classification dataset would likely improve the effectiveness of theme scoping and, consequently, the overall efficiency gains it provides.
@@
These limitations suggest several promising directions for future research, including improved knowledge extraction techniques, more robust theme classification approaches, and enhanced graph evolution strategies. A particularly promising direction is the integration of external knowledge graphs into RAS’s iterative loop, as its graph-based architecture naturally accommodates the incorporation of established KGs like Wikidata or domain-specific knowledge bases. This could significantly enhance the system’s reasoning capabilities by combining dynamically constructed knowledge with curated structured information. We summarize the potential future directions of extension in Figure 5.
@@ Appendix B HotpotQA-SUBQ Dataset
The source of our dataset is HotpotQA (Yang et al., 2018).222We use the version hotpot_train_v1.1 from https://github.com/hotpotqa/hotpot. Our goal is to create a dataset that can let the model learn the capabilities of subquery generation and answering with intensive knowledge. We name the dataset as HotpotQA-SUBQ.
@@ B.3 Sample Labeling
The sample labeling process transforms the filtered and subquery-augmented HotpotQA examples into training instances for both the Action Planner and Answerer components. We formalize this process in Algorithm 1.
The algorithm takes as input the filtered HotpotQA dataset , base language model , and text-to-triple conversion model . For each example, we first attempt direct answer generation using the base LLM (Line 3). If successful, we create [NO_RETRIEVAL] training instances for both the planner and answerer (Lines 4-6).
@@
Label planner sample as [SUFFICIENT]
Create answerer training instance with ground truth answer
@@
: Trains the Action Planner to determine retrieval needs and generate targeted subqueries
: Trains the Answerer to synthesize final responses from accumulated graph context
Table 3 illustrates the distribution of our generated datasets. The planner dataset demonstrates a balanced distribution across three label types: 35% [SUFFICIENT], 55% [SUBQ], and 10% [NO_RETRIEVAL]. The answerer dataset incorporates both no-retrieval and sufficient cases from multiple sources. The size of the answerer dataset exceeds the combined total of no-retrieval and sufficient cases from the planner dataset for two reasons: first, we incorporated additional data from ASQA and Arc-Easy datasets; second, we included no-retrieval cases that were initially filtered out in step 1 of our process. This comprehensive approach ensures a robust training set for the answerer component.
@@ B.4 Dataset Statistics
| Planner Data | Answerer Data | ||
@@ Appendix C Training Details
C.1 Classifier & Shifter Training for Theme Scoping
Classifier Training. For theme classification, we trained a multilabel document classifier on the DBPedia-298 dataset (Lehmann et al., 2015) using a two-stage architecture. The first stage uses the nomic-embed-text-v1 encoder model (Nussbaum et al., 2024) to generate document embeddings, followed by a classification head consisting of a two-layer neural network. The training size and the validation size are 240,932 and 36,008, respectively.
The classifier consists of:
An encoder based on nomic-embed-text-v1 for generating document embeddings
A classification head with:
Input layer: embedding dimension → 512 units
ReLU activation
Dropout (p=0.1)
Output layer: 512 → number of labels (298)
Sigmoid activation for multilabel outputs
We jointly trained both the encoder and classification head with different learning rates:
Classification head: Trained with a higher learning rate (2e-4) to allow faster adaptation to the classification task
Text Encoder: Fine-tuned with a lower learning rate (2e-5) to preserve pretrained knowledge while making subtle adjustments for the domain
This dual learning rate strategy allows the model to benefit from the pretrained knowledge in the encoder while optimizing the classification head for the specific multilabel task. The lower learning rate for the encoder helps prevent catastrophic forgetting of the pretrained representations.
The model was evaluated using micro-averaged metrics on the validation set during training, with monitoring of F1 score, precision, and recall, as shown in Figure 16. We select the best classifier model based on the best validation F1 score achieved during training.
Distribution Shifter Training. The theme distribution shifter maps query theme distributions to expected document theme distributions. Training involves two phases:
Phase 1: PPD Labeling
Using our constructed HotpotQA-SUBQ dataset (in Appendix B) with 204,638 query-doc pairs (where we set training-validation ratio as 9:1):
Input Generation: Process all subqueries and supporting documents through the trained theme classifier to obtain their Posterior Probability Distributions (PPDs)
Pairing: Each subquery is paired with its corresponding supporting document, creating query-document PPD pairs
Data Structure: The labeled dataset consists of:
Input: Query PPDs (298-dimensional probability vectors)
Target: Corresponding document PPDs (298-dimensional probability vectors)
Phase 2: Shifter Training
Train a neural network to map query PPDs to document PPDs:
Architecture: A three-layer MLP with:
Input layer: 298 → 512 units
Hidden layer: 512 → 256 units
Output layer: 256 → 298 units
ReLU activations and dropout (p=0.2) between layers
Softmax activation for final output
Training Objectives: Multiple distribution alignment metrics:
KL divergence loss (primary objective)
Jensen-Shannon divergence monitoring
Wasserstein distance monitoring
L1 and L2 distance monitoring
Training Configuration:
Optimizer: Adam with learning rate 1e-3
Batch size: 32
Early stopping based on validation loss
Numerical stability ensured through epsilon addition (1e-15)
Performance is tracked using validation KL divergence, Jensen-Shannon divergence, and Wasserstein distance to ensure proper distribution alignment, as shown in Figure 16.
C.2 Text-to-Triples Model Training
For our text-to-triple conversion model, we fine-tuned a Flan-T5-Large model (Chung et al., 2024) to transform raw text passages into structured knowledge triples. The model processes input text sequences up to 512 tokens in length and generates structured triples in a standardized format “(S subject P predicate O object)”. 333Our fine-tuned can be found at https://huggingface.co/pat-jj/text2triple-flan-t5 and https://huggingface.co/pat-jj/text2graph-llama-3.2-3b
Our training dataset WikiOFGraph (Kim et al., 2024a), curated by ChatGPT, comprises 5,851,776 text-triple pairs, with an additional 5,000 samples reserved for validation. Each training instance consists of a natural language text passage paired with its corresponding set of comma-separated triples. For example:
@@ C.2 Text-to-Triples Model Training
C.3 Action Planner & Answerer Multi-task Training
We employ GraphLLM (Perozzi et al., 2024; He et al., 2024) as our foundational model architecture and utilize our constructed HotpotQA-SUBQ dataset to train a unified model capable of performing both graph-conditioned action planning and answer generation through multitask learning.
Since the Action Planner and Answerer share an identical architectural foundation, as illustrated in Figure 1, we differentiate their functionality through specialized instruction sets. For the Action Planner , we employ the following instruction template:
For the Answerer , we utilize this distinct instruction set:
For graph encoding, we implement Graph Transformer (Shi et al., 2020), selected for its robust capability in handling non-fully-connected graphs—a common characteristic of text-extracted triples. Our base language models comprise LLaMA-2-7B and LLaMA-3-8B, chosen both to maintain consistency with previous research (Asai et al., 2023; Lyu et al., 2024b) and to investigate our framework’s performance scaling across different model capacities.
Our implementation of GraphLLM differs from G-Retriever (He et al., 2024) primarily due to the distinct nature of our graph structures. While G-Retriever operates on single interconnected graphs, our framework processes multiple potentially disconnected subgraphs, each corresponding to different subqueries. To address this architectural difference, we adopt a sequential encoding strategy: rather than encoding the entire graph at once, we process each subgraph individually using Graph Transformer, followed by mean pooling across all subgraph embeddings to produce the final encoded representation.
@@ D.1 Test Datasets
We evaluate RAS on diverse benchmark datasets spanning short-form QA, closed-set tasks, and long-form generation in the zero-shot setting, aligning with (Asai et al., 2023; Lyu et al., 2024b). Below we describe each dataset:
@@
TriviaQA-unfiltered (TQA) (Joshi et al., 2017): A large-scale QA dataset containing 11,313 question-answer pairs in our test set. The questions are sourced from trivia enthusiasts and cover diverse topics.
2WikiMultiHopQA (2WQA) (Ho et al., 2020): A multi-hop question answering dataset (with 12,576 samples in test set) that requires models to combine information from multiple Wikipedia articles to answer questions.
PopQA (Mallen et al., 2022): A dataset focusing on questions about long-tail entities, containing 1,399 queries where the monthly Wikipedia page views are less than 100. These questions test models’ ability to handle queries about less common entities.
@@
PubHealth (Pub) (Zhang et al., 2023a): A dataset for verifying health-related claims. Models must classify statements as ”true” or ”false” based on scientific evidence. The dataset contains 987 test samples.
ARC-Challenge (ARC) (Clark et al., 2018): A multiple-choice science question dataset designed to be challenging for models, requiring multi-hop reasoning and background knowledge. The dataset contains 1,172 test samples.
@@
ASQA (Stelmakh et al., 2022): A long-form question answering dataset that requires models to generate comprehensive answers with proper citation of sources. Models must balance information completeness with factual accuracy. The dataset contains 948 test samples.
ELI5 (Fan et al., 2019): A dataset derived from the ”Explain Like I’m Five” subreddit, containing questions seeking straightforward explanations of complex topics. The dataset contains 1,000 test samples.
@@ D.2 Evaluation Metrics
For PopQA and TriviaQA, we use the ”golden match” metric (Asai et al., 2023; Min et al., 2019; Guu et al., 2020)444Implementation aligned with https://github.com/AkariAsai/self-rag/blob/main/retrieval_lm/metrics.py, where a prediction is considered correct if it contains any normalized version of the ground truth answers :
@@
For 2WikiMultiHopQA, we follow RPG (Lyu et al., 2024b)555Implementation aligned with https://github.com/haruhi-sudo/RPG/blob/main/retriever/src/evaluation.py to use token-level F1 score between prediction and ground truth :
@@
Closed-set Tasks: For both PubHealth and ARC-Challenge, we use accuracy (Asai et al., 2023), computed as:
@@ E.1 Closed-Source Model Settings
For Action Planning (Figure 21), we design a comprehensive prompt that guides the model in making strategic decisions about information retrieval needs. The prompt instructs the model to output one of three labels based on careful analysis of the current knowledge state: (1) [NO_RETRIEVAL] when the question can be answered directly, either due to the model’s inherent knowledge or the question’s nature, (2) [SUBQ] accompanied by a focused subquery when additional information is needed, or (3) [SUFFICIENT] when the accumulated knowledge is adequate to answer the question. The prompt includes diverse examples demonstrating different scenarios:
@@
For Answer Generation (Figure 22), we design a prompt that focuses on synthesizing precise answers from structured knowledge graphs. The prompt emphasizes selective use of retrieved information - instructing the model to utilize subqueries and graph information only when relevant to the question at hand. A distinctive feature of this prompt is its requirement for definitive answers even under uncertainty or incomplete information, ensuring the model always provides a response.
@@
In addition, for ASQA and ELI5 with closed-source models (RAS and all the other baselines (e.g., SuRe) using Sonnet-3.5), we conduct few-shot inference with two in-context learning demonstrations, aligning with previous study’s (Gao et al., 2023) implementation.777https://github.com/princeton-nlp/ALCE/blob/main/run.py888Exemplars can be found at https://github.com/princeton-nlp/ALCE/tree/main/prompts
@@ E.2 Open-Source Model Settings
For more efficient text retrieval, we split both theme index and dense index of the corpus into eight splits, and load them on eight GPUs with faiss-gpu 1.7.2.999All the Wikipedia dumps (2018, 2020) we used are downloaded from https://github.com/facebookresearch/atlas
We set maximum new tokens as 100 for PopQA, TriviaQA, and 2WikiMultihopQA, as 50 for PubHealth and ARC-Challenge, and as 300 for ASQA and ELI5, aligning with previous study (Asai et al., 2023; Lyu et al., 2024b). All the generation are configured with batch size of 1.
As we always set the main query as the first subquery, we pre-extract the triples from the context at the first iteration using our Flan-T5-based text-to-triple model for batch triple extraction. Notably, when we experimented with using Claude 3.5 Sonnet for triple extraction using the prompt from Figure 20, we observed an additional 3% average performance gain. This suggests room for improvement in our triple extraction approach - one potential direction would be training a larger model like LLaMA-2-7B on the WikiOFGraph dataset to achieve higher-quality triple extraction.
@@ E.3 Instructions
Appendix F Qualitative Examples

### v2

@@ 5 Conclusion
We presented RAS (Retrieval-And-Structuring), a novel framework that improves knowledge-intensive language generation by interleaving iterative retrieval with dynamic knowledge graph construction. RAS introduces key innovations, including structure-aware action planning and graph-augmented answering, enabling models to reason over progressively organized information tailored to each query. Extensive experiments across seven benchmarks demonstrate that RAS consistently outperforms existing methods, particularly in tasks requiring multi-step reasoning and factual consistency. Furthermore, the modular, graph-based architecture of RAS allows seamless integration with external knowledge sources and human feedback, making it well-suited for high-stakes domains where transparency and adaptability are essential. Limitations and directions for future work are discussed in Appendix A.
@@ Appendix A Limitations & Future Work
Knowledge Extraction Quality: The effectiveness of both planning and answer generation depends on the quality of text-to-triple extraction. Despite using an end-to-end text-to-triples model for triple extraction, implementing a more sophisticated approach – such as a dedicated Named Entity Recognition and Relation Extraction (NER-RE) pipeline – could potentially enhance RAS’s reasoning capabilities and knowledge representation.
@@
These limitations suggest several promising directions for future research, including improved knowledge extraction techniques and enhanced graph evolution strategies. A particularly promising direction is the integration of external knowledge graphs into RAS’s iterative loop, as its graph-based architecture naturally accommodates the incorporation of established KGs like Wikidata or domain-specific knowledge bases. This could significantly enhance the system’s reasoning capabilities by combining dynamically constructed knowledge with curated structured information. We summarize the potential future directions of extension in Figure 5.
@@ Appendix B HotpotQA-SUBQ Dataset
The source of our dataset is HotpotQA [1].111We use the version hotpot_train_v1.1 from https://github.com/hotpotqa/hotpot. Our goal is to create a dataset that can let the model learn the capabilities of subquery generation and answering with intensive knowledge. We name the dataset as HotpotQA-SUBQ.
@@ B.3 Sample Labeling
The sample labeling process transforms the filtered and subquery-augmented HotpotQA examples into training instances for both the Planning and Answering components. We formalize this process in Algorithm 1.
The algorithm takes as input the filtered HotpotQA dataset , base language model , and text-to-triple conversion model . For each example, we first attempt direct answer generation using the base LLM (Line 3). If successful, we create [NO_RETRIEVAL] training instances for both the planning and answering (Lines 4-6).
@@
Label planning sample as [SUFFICIENT]
Create answering training instance with ground truth answer
@@
: Trains the planning ability to determine retrieval needs and generate targeted subqueries
: Trains the answering ability to synthesize final responses from accumulated graph context
Table 4 illustrates the distribution of our generated datasets. The planning dataset demonstrates a balanced distribution across three label types: 35% [SUFFICIENT], 55% [SUBQ], and 10% [NO_RETRIEVAL]. The answering dataset incorporates both no-retrieval and sufficient cases from multiple sources. The size of the answering dataset exceeds the combined total of no-retrieval and sufficient cases from the planning dataset for two reasons: first, we incorporated additional data from ASQA and Arc-Easy datasets; second, we included no-retrieval cases that were initially filtered out in step 1 of our process. This comprehensive approach ensures a robust training set for the answering component.
@@ B.4 Dataset Statistics
| Planning Data | Answering Data | ||
@@ Appendix C Training Details
C.1 Text-to-Triples Model Training
For our text-to-triple conversion model, we fine-tuned a Flan-T5-Large model [69]222https://huggingface.co/pat-jj/text2triple-flan-t5 and a LLaMA-3.2-3B-Instruct model [70]333https://huggingface.co/pat-jj/text2graph-llama-3.2-3b to transform raw text passages into structured knowledge triples. The model processes input text sequences up to 512 tokens in length and generates structured triples in a standardized format “(S subject P predicate O object)”.
Our training dataset WikiOFGraph [44], curated by ChatGPT, comprises 5,851,776 text-triple pairs, with an additional 5,000 samples reserved for validation. Each training instance consists of a natural language text passage paired with its corresponding set of comma-separated triples. For example:
@@ C.1 Text-to-Triples Model Training
C.2 Multitask Training of Planning and Answering Capabilities
We employ Graph LLM [38, 11] as our foundational model architecture and utilize our constructed HotpotQA-SUBQ dataset to train a unified model capable of performing both graph-conditioned action planning and answer generation through multitask learning.
Since both tasks share an identical architectural foundation, as illustrated in Figure 1, we differentiate their functionality through specialized instruction sets. For Planning, we employ the following instruction template:
For the Answering, we utilize this distinct instruction set:
For graph encoding, we implement Graph Transformer [65], selected for its robust capability in handling non-fully-connected graphs—a common characteristic of text-extracted triples. Our base language models comprise LLaMA-2-7B and LLaMA-3-8B, chosen both to maintain consistency with previous research [30, 31] and to investigate our framework’s performance scaling across different model capacities.
Our implementation of GraphLLM differs from G-Retriever [11] primarily due to the distinct nature of our graph structures. While G-Retriever operates on single interconnected graphs, our framework processes multiple potentially disconnected subgraphs, each corresponding to different subqueries. To address this architectural difference, we adopt a sequential encoding strategy: rather than encoding the entire graph at once, we process each subgraph individually using Graph Transformer, followed by mean pooling across all subgraph embeddings to produce the final encoded representation.
@@ D.1 Test Datasets
We evaluate RAS on diverse benchmark datasets spanning short-form QA, closed-set tasks, and long-form generation in the zero-shot setting, aligning with [30, 31]. Below we describe each dataset:
@@
TriviaQA-unfiltered (TQA) [51]: A large-scale QA dataset containing 11,313 question-answer pairs in our test set. The questions are sourced from trivia enthusiasts and cover diverse topics.
2WikiMultiHopQA (2WQA) [53]: A multi-hop question answering dataset (with 12,576 samples in test set) that requires models to combine information from multiple Wikipedia articles to answer questions.
PopQA [52]: A dataset focusing on questions about long-tail entities, containing 1,399 queries where the monthly Wikipedia page views are less than 100. These questions test models’ ability to handle queries about less common entities.
@@
PubHealth (Pub) [54]: A dataset for verifying health-related claims. Models must classify statements as "true" or "false" based on scientific evidence. The dataset contains 987 test samples.
ARC-Challenge (ARC) [55]: A multiple-choice science question dataset designed to be challenging for models, requiring multi-hop reasoning and background knowledge. The dataset contains 1,172 test samples.
@@
ASQA [57]: A long-form question answering dataset that requires models to generate comprehensive answers with proper citation of sources. Models must balance information completeness with factual accuracy. The dataset contains 948 test samples.
ELI5 [58]: A dataset derived from the "Explain Like I’m Five" subreddit, containing questions seeking straightforward explanations of complex topics. The dataset contains 1,000 test samples.
@@ D.2 Evaluation Metrics
For PopQA and TriviaQA, we use the "golden match" metric [30, 71, 8]444Implementation aligned with https://github.com/AkariAsai/self-rag/blob/main/retrieval_lm/metrics.py, where a prediction is considered correct if it contains any normalized version of the ground truth answers :
@@
For 2WikiMultiHopQA, we follow RPG [31]555Implementation aligned with https://github.com/haruhi-sudo/RPG/blob/main/retriever/src/evaluation.py to use token-level F1 score between prediction and ground truth :
@@
Closed-set Tasks: For both PubHealth and ARC-Challenge, we use accuracy [30], computed as:
@@ E.1 Closed-Source Model Settings
For Planning (Figure 19), we design a comprehensive prompt that guides the model in making strategic decisions about information retrieval needs. The prompt instructs the model to output one of three labels based on careful analysis of the current knowledge state: (1) [NO_RETRIEVAL] when the question can be answered directly, either due to the model’s inherent knowledge or the question’s nature, (2) [SUBQ] accompanied by a focused subquery when additional information is needed, or (3) [SUFFICIENT] when the accumulated knowledge is adequate to answer the question. The prompt includes diverse examples demonstrating different scenarios:
@@
For Answering (Figure 20), we design a prompt that focuses on synthesizing precise answers from structured knowledge graphs. The prompt emphasizes selective use of retrieved information - instructing the model to utilize subqueries and graph information only when relevant to the question at hand. A distinctive feature of this prompt is its requirement for definitive answers even under uncertainty or incomplete information, ensuring the model always provides a response.
@@
In addition, for ASQA and ELI5 with closed-source models (RAS and all the other baselines (e.g., SuRe) using Sonnet-3.5), we conduct few-shot inference with two in-context learning demonstrations, aligning with previous study’s [56] implementation.777https://github.com/princeton-nlp/ALCE/blob/main/run.py888Exemplars can be found at https://github.com/princeton-nlp/ALCE/tree/main/prompts
@@ E.2 Open-Source Model Settings
For more efficient text retrieval, we split the dense index of the corpus into eight splits, and load them on eight GPUs with faiss-gpu 1.7.2.999All the Wikipedia dumps (2018, 2020) we used are downloaded from https://github.com/facebookresearch/atlas
We set maximum new tokens as 100 for PopQA, TriviaQA, and 2WikiMultihopQA, as 50 for PubHealth and ARC-Challenge, and as 300 for ASQA and ELI5, aligning with previous study [30, 31]. All the generation are configured with batch size of 1.
As we always set the main query as the first subquery, we pre-extract the triples from the context at the first iteration using our text-to-triple model for batch triple extraction.
@@ E.3 Instructions
Appendix F Supplementary Results and Analysis
Higher Performance with More Contexts?
We further examine the performance of Sonnet-3.5 when increasing the number of retrieved documents. Table 6 compares Sonnet-3.5 under different retrieval depths with RAS-Sonnet.
| Method | 2WikiMultihopQA (F1) | PopQA (Acc) | PubHealth (Acc) | ARC (Acc) |
| Sonnet-3.5 (#docs=5) | 53.7 | 57.3 | 53.9 | 87.1 |
| Sonnet-3.5 (#docs=10) | 55.1 | 58.5 | 56.6 | 87.9 |
| Sonnet-3.5 (#docs=15) | 55.4 | 58.0 | 56.9 | 86.9 |
| RAS (Sonnet) | 57.7 | 62.3 | 71.4 | 93.8 |
Increasing retrieval depth yields diminishing returns for Sonnet-3.5, while RAS-Sonnet achieves significantly higher performance with structured, targeted context. This highlights the limitations of passive context scaling and the benefits of RAS’s organized retrieval strategy.
Error Analysis on Graph Construction
We present two representative cases where the quality of retrieved structured knowledge significantly impacts final answer generation. These examples are distinct from the qualitative examples shown in Section G and highlight both successful and failed reasoning paths.
Case 1: Accurate Reasoning with Well-Structured Graph
Question: Who directed the film adaptation of the novel “Tinker Tailor Soldier Spy”?
Subqueries and Graphs:
[SUBQ] Who directed the film “Tinker Tailor Soldier Spy”?
(S> Tinker Tailor Soldier Spy | P> Film adaptation director | O> Tomas Alfredson),
(S> Tinker Tailor Soldier Spy | P> Release year | O> 2011) -
[SUBQ] Who is Tomas Alfredson?
(S> Tomas Alfredson | P> Occupation | O> Film director)
Answer: Tomas Alfredson (Correct)
Analysis: The subquery identifies the correct adaptation and its director. The structured triple precisely encodes this relation, allowing the model to reason accurately.
Case 2: Reasoning Failure Due to Spurious or Missing Triples
Question: What invention is Nikola Tesla best known for?
Subqueries and Graphs:
[SUBQ] What are Nikola Tesla’s major inventions?
(S> Nikola Tesla | P> Known for | O> Earthquake machine),
(S> Nikola Tesla | P> Proposed | O> Death ray),
(S> Nikola Tesla | P> Created | O> Tesla coil) -
[SUBQ] What is a Tesla coil used for?
(S> Tesla coil | P> Used in | O> High-frequency experiments)
Answer: Earthquake machine or death ray (Incorrect)
Analysis: The graph omits the key triple (Nikola Tesla | Invented | Alternating current system), instead surfacing fringe inventions. This misleads the model into emphasizing less impactful or speculative work, leading to an incorrect answer.
These examples reveal that RAS makes reasoning pathways more interpretable by structuring retrieved knowledge explicitly. When failures occur, they can often be traced to identifiable issues in the extracted triples, such as missing key relations or overemphasis on fringe facts, rather than opaque model behavior. This transparency not only facilitates error analysis but also suggests concrete avenues for improving upstream components like triple extraction or subquery generation, making RAS a flexible foundation for future refinement.
Appendix G Qualitative Examples